OpenStudy (anonymous):

evaluate g(n-5) if g(x)= x^2-6 / 2x Please help me with this multiple choices. A. n^2 - 5 / 2n-20 B. n^2-6/2n -5 C. n^2-10n+19 / 2n-10

hero (hero):

@carlos10, rather than close the previous question, you can just bump the question. I would have shown you how to simplify it.

OpenStudy (anonymous):

to choose one of the answers!

OpenStudy (anonymous):

please!

OpenStudy (anonymous):

think of the function like this... \[g()=()^{2} -\frac{ 6 }{ 2() }\] so whatever is in the parentheses on the left goes in th parentheses on the right. \[g(n-5)=(n-5)^{2} -\frac{ 6 }{ 2(n-5) }\] now just simplify...

hero (hero):

@pgpilot326, I'm afraid you have interpreted the function too literally. It is \[g(n - 5) = \frac{(n - 5)^2 - 6}{2(n - 5)}\] Right @carlos10 ?

OpenStudy (anonymous):

oh, sorry. it's difficult to tell when people don't use parentheses at all or don't use them properly.

hero (hero):

I agree. It's too bad that people don't know how to post equations horizontally.

OpenStudy (anonymous):

yes it is but can someone teach me how to simplify it?

OpenStudy (anonymous):

please.

OpenStudy (anonymous):

you got this @Hero ?

OpenStudy (anonymous):

to get the right answer.

OpenStudy (anonymous):

I want to learn how to simplify it really fast please

hero (hero):

I was ready to help until I figured out someone was taking a test. Then I got discouraged very quickly.

OpenStudy (anonymous):

Im done with that, the test is over but now I want to learn how can I simplify it!